OpenStudy (anonymous):

Given, \[{x_1} > a > 0,\] \[{x_{n + 1}} = \sqrt {{x_n}^2 - 2a{x_n} + 2{a^2},} \] \[n = 1,2,...\] Prove that \[\mathop {\lim }\limits_{n \to \infty } {x_n}\] is convergent and compute its value.

OpenStudy (jamesj):

Suppose for a moment it is convergent and has limit L, then \[ L = \sqrt{L^2 - 2aL + 2a^2} \] \[ \implies L^2 = L^2 - 2aL + 2a^2 \] \[ \implies L = a \] Now, to prove the sequence actually is convergent, show that (xn) is monotonically decreasing and bounded below.

OpenStudy (anonymous):

That is great, but can you show me more detail about why xn is decreasing?

OpenStudy (jamesj):

That's a challenge for you to show. I recommend trying to show in general for x > a that \[ \sqrt{x^2 - 2ax + 2a^2} \leq x \]

OpenStudy (jamesj):

This statement is true iff \[ x^2 - 2ax + 2a^2 \leq x^2 \] which is true iff \[ a \leq x \]

OpenStudy (jamesj):

So to show the sequence is monotonically decreasing, you actually want to show it is bounded below by a.

OpenStudy (jamesj):

So now if show that, then you will have that it is indeed bounded below and by the calculation above, it is also monotonically decreasing.

OpenStudy (anonymous):

I can see \[{x_2}^2 - {x_1}^2 = 2a(a - {x_1}) < 0\] \[{x_3}^2 - {x_2}^2 = 2a(a - {x_2}) < 0?\]

OpenStudy (anonymous):

and why all that become \[{x^2} - 2ax + 2{a^2} \leqslant {x^2}\]

OpenStudy (jamesj):

If we show for ANY x > a, that \[ \sqrt{x^2 - 2ax + 2a^2} < x \] then it will follow that \[ x_{n+1} < x_n \ \ \hbox{ for all } n\] Now \[ \sqrt{x^2 - 2ax + 2a^2} < x \] is true if and only if, squaring both sides, and as both sides are positive the inequality is preserved, \[ x^2 - 2ax + 2a^2 < x^2 \] i.e., \[ 2a^2 < 2ax \] i.e., \[ a < x \] Hence if we can show that \( x_n > a \) for all n, then it follows that the sequence \( (x_n) \) is also monotonically decreasing.

OpenStudy (anonymous):

I know x1>a , But why xn > a ?

OpenStudy (jamesj):

To show that is true start with an arbitrary number x > a. Show that that implies \[ \sqrt{x^2 - 2ax + 2a^2} > a --- (*)\] That being true, it follows by induction that \( x_n > a \) for all n. (And that is because we know that \( x_1 > a \) and the relationship (*) shows that \( x_k > a \implies x_{k+1} > a \) for arbitrary k)

OpenStudy (anonymous):

I belive you have try your best to explain the reason. But ..I don't get it..Let me see more book. ..I wish you are my classmate.

OpenStudy (jamesj):

ok. To recap the overall method here: - show (xn) is bounded below - show (xn) is decreasing - then by a standard and very important theorem in analysis, (xn) is a convergent series - knowing now that (xn) is convergent, write L for it's limit; we can solve for L and find it is a.

OpenStudy (jamesj):

Have a look at your lecture notes or text book. I hope/trust there is another example there going through these sorts of steps.

OpenStudy (anonymous):

En. I will. Thank you so much.

OpenStudy (anonymous):

I got it, JamesJ . haha. Thanks. If I ask other question, please help me, and I promise my question will be more interesting.